OpenStudy (anonymous):

how do i factor sin^2x csx^2x - sin^2x

OpenStudy (anonymous):

csc*

OpenStudy (anonymous):

Take \(sin^2 x\) common from both the terms..

OpenStudy (anonymous):

\[\huge = \sin^2x(cosec^2x - 1)\]

OpenStudy (valpey):

Um, isn't \[sin^2(x)csc^2(x) = \frac{sin^2(x)}{sin^2(x)} = 1\]

OpenStudy (valpey):

And isn't \[1-sin^2(x) = cos^2(x)\]

OpenStudy (anonymous):

He is asking to factorize and not to solve I think..

OpenStudy (anonymous):

if i simplified the expression would i get -1? o-o

OpenStudy (valpey):

if you simplify the expression you get \[cos^2(x)\]

OpenStudy (anonymous):

using fundamental identities

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

after i factor it and simplify

OpenStudy (valpey):

yes \[sin^2(x)csc^2(x)-sin^2(x) = \frac{sin^2(x)}{sin^2(x)}-sin^2(x) =1-sin^2(x)=cos^2(x)\]

OpenStudy (valpey):

Just need definition of csc and the pythagorean identity: \[sin^2(x)+cos^2(x)=1\]

OpenStudy (anonymous):

but when i factor i get sin^2x (csc^2x-1) ?

OpenStudy (valpey):

That is equivalent, but irrelevant.

OpenStudy (valpey):

csc = 1/sin

OpenStudy (anonymous):

sin^2x(cosec^2x-1) .................... since: cosecx=1/sinx therefore= sin^2x(1/sin^2x-1) multiple sin^2x therefore=sin^4x(1-sin^2x).............................since:(sin^2x+cos^x)=1 therefore=sin^4x(cos^2x+sin^2x-sin^2x) therefore=sin^4xcos^2x=sin^2x(sin^2xcos^2x) since:(2sinxcosx)=sin2x therefore=sin^2x[0.5sin^2(2x)]

OpenStudy (anonymous):

I HOPE IT WORKS !!

OpenStudy (valpey):

sin = O/H cos = A/H tan = O/A csc = H/O sec = H/A cot = A/O What you just asked is the same as: \[(O/H)^2(H/O)^2-(O/H)^2\]

OpenStudy (anonymous):

i did i/csc^2x (csc^2x-1) can i cancel out the csc^2x?

OpenStudy (valpey):

\[\frac{O^2}{H^2}\frac{H^2}{O^2}−\frac{O^2}{H^2} =1-\frac{O^2}{H^2}=\frac{H^2}{H^2}-\frac{O^2}{H^2}=\frac{H^2-O^2}{H^2}\]

OpenStudy (valpey):

But Pythagorus taught us: \[H^2=O^2+A^2\ \ or\ \ H^2-O^2=A^2 \] \[\frac{H^2-O^2}{H^2}=\frac{A^2}{H^2}= (\frac{A}{H})^2 = cos^2(x)\]

OpenStudy (valpey):

When in doubt, SOHCAHTOA!!!!

OpenStudy (anonymous):

haha, i was confused by the variables at first but alright, thanks!

OpenStudy (valpey):

The fact that the trigonometric relationships are defined as these ratios and their relationship in a right triangle is fundamental. Make sure you understand this aspect of trigonometry if you commit anything to memory.